Practice Exercise 4.6





Pure aluminum (Al) metal reacts with hydrochloric acid (HCI) to produce aluminum chloride (AlCl3) and hydrogen (H2) gas.





a. Write the balanced chemical equation for the reaction described above.




b. If the expected mass of AlCl3 to be formed is 3.5 g, what should be the initial mass of the pure aluminum metal?




c. If the actual mass of AlCl3 produced is 3.1 g, what is the percentage yield of the reaction?





1
Expert's answer
2021-11-22T02:58:01-0500

a) 2Al+6HCl→2AlCl3+3H2 2 A l + 6 H C l → 2 A l C l 3 + 3 H 2 .

b) B. 2Al + 3Cl2 = 2AlCl3

2AlCl3 = 2Al

2 × 133g AlCl3 = 2 ×27 g Al

1g AlCl3 = 2×27/(2×133) g Al

3.5g AlCl3 = 27 × 3.5 g/ 133 Al

= 0.71g Al

C. % yield = actual yield / theoretical yield × 100%

= 3.1/3.5 ×100%

= 88.57%